OpenStudy (lisa):

The equation 9(u – 2) + 1.5u = 8.25 models the total miles Michael traveled one afternoon while sledding, where u equals the number of hours walking up a hill and (u – 2) equals the number of hours sledding down the hill. Which is the value of u? u = 0.25 u = 0.75 u = 1.1 u = 2.5

OpenStudy (anil):

9(u-2) +1.5 u =8.25 9u- 9*2+1.5u=8.25 10.5 u -18=8.25 10.5 u = 8.25+18 10.5u =26.25 u=26.25/10.5 u=2.5